OpenStudy (anonymous):

ln(x/((x^2)-1)) differentiate the function

OpenStudy (anonymous):

\[\huge \ln(\frac{x}{x^2-1}) ?\]

OpenStudy (anonymous):

use substitution: \[\huge \frac{d}{dx}\ln(u)=\frac{1}{u}*\frac{du}{dx}\] where by the derivative of the inner function must be performed by the quotient rule